OpenStudy (anonymous):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line .y=7/5x-3/5

OpenStudy (anonymous):

A. 5x + 7y = 1 B. 7x + 5y = –13 C. 7x + 5y = –1 D. 3x + 5y = 7

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."